OpenStudy (anonymous):

If the point (a,3) lies on the graph of the equation 5x + y = 8, then a=

OpenStudy (anonymous):

(a,3) means that when y=3, x=a. So, we plug y=3 into the equation, and solve it for x: 5x + y = 8 5x + 3 = 8 (because y=3) 5x = 5 (8-3) x = 1 (5÷5) a = 1 (Because x = a)

OpenStudy (anonymous):

Indicate the equation of the line through (2, -4) and having slope of 3/5.

OpenStudy (anonymous):

Do you know the formula for point-gradient form? y-y₁ = m(x-x₁) y+4 = 3/5(x-2)

OpenStudy (anonymous):

No I dont Actually.

OpenStudy (anonymous):

It's y-y₁ = m(x-x₁), where x₁ and y₁ are the x and y co-ordinates of the point you're given, and m is the slope of the line.

OpenStudy (anonymous):

Ok Thanks alot.